OpenStudy (anonymous):

What is the radius of a circle with a circumference of twenty-six point six nine meters? Use three point one four for pi and show your work. @hello_j

OpenStudy (anonymous):

2*pi*r=26.69

OpenStudy (anonymous):

The formula for the circumference is 2 pi radius. All you have to do is plug 26.69 and try to isolate the radius for your answer :)

OpenStudy (anonymous):

r=26.69/2*3.14=

OpenStudy (anonymous):

r=3.903

OpenStudy (anonymous):

So I write r=26.69/2*3.14= 3.903 r=3.903?

OpenStudy (anonymous):

@dg2

OpenStudy (anonymous):

sorry r=4.25 .

OpenStudy (anonymous):

now okay?